Melinda's father Lester is 3 times as old as Melinda. Seven years ago, he was 4 times as old as Melinda. How old is each now?
Answer by stanbon(75887)

Melinda's father Lester is 3 times as old as Melinda. Seven years ago, he was 4 times as old as Melinda. How old is each now?
------
Equations:
f = 3m
(f-7) = 4(m-7)
------
Substitute for "f" and solve for "m":
(3m-7) = 4m-28
m = 21 (Melinda's age now)
=============
Solve for "f":
f = 3m
f = 3*21 = 63 (Father's age now)
===================================
cheers,
Stan H.